# The Science of Data Collection: Insights from Surveys can Improve Machine Learning Models

### v1

@@ 3.3 Mitigation Measures and Future Research
The survey methods literature suggests several approaches to minimize the effects introduced in Section 3.2.
@@
Many surveys spend weeks or even months drafting, testing, and revising questions and response options to arrive at language that is understood similarly by most members of the population, a process called cognitive interviewing (Willis, 2004). They then launch the survey with a small group of respondents to assess response time, don’t know rates, and response rates. Such testing could improve the instructions and prompts given to labelers.
@@
Many surveys capture process data, called paradata, during the survey, such as the time spent on each screen, the device used, even mouse movements (Kreuter, 2013; Horwitz et al., 2017). Paradata can help identify satisficing respondents and low quality data. While paradata allows for additional insights, collecting such data may raise additional privacy and ethics concerns. For a discussion of the privacy and ethics of paradata collection in surveys, see Couper & Singer (2012); Kunz et al. (2020); Henninger et al. (2023).
@@
Label collection instruments could experiment with prompts to encourage labelers not to engage in satisficing. Respondents who pick many ‘don’t know’ answers or repeatedly choose the same response option can receive reminders about the importance of the task. Those who click through screens more quickly than they could possibly read questions can receive prompts to slow down and read carefully. In surveys, feedback on speeding successfully slowed respondents and did not lead to early terminations of the survey (Conrad et al., 2017).
Limit observations per labeler
Many labeling tasks allow labelers to continue as long as they like or until there are no more observations to label. However, satisficing may be more likely as the task continues, as found in surveys (Galesic et al., 2008). Thus, ML and AI researchers may wish to limit labelers to a maximum number of observations. However, such a limit may not make sense for tasks with high training costs.
@@
In surveys, instructed response items, such as “Choose yellow below” can help identify respondents who speed or provide low quality responses (Gummer et al., 2018). Embedding observations with known labels in a labeling task can similarly help identify satisficing labelers (Nie et al., 2020). In NLP tasks, it is common to qualify only workers who pass an initial quiz or perform well on inserted test observations Nangia et al. (2021).
@@
We recommend future work to more comprehensively test these and related approaches.
4 Labelers as Interviewers
Thinking of labelers as interviewers rather than respondents leads to additional insights.
4.1 Interviewer effects in surveys
When survey data are collected by interviewers, as in face-to-face or phone surveys, the characteristics and behavior of the interviewer, and the interaction of these with the respondents and the questions, impact the data. For example, interviewer age, gender, and race can influence respondents’ answers (see West & Blom, 2016, for a detailed review).
The statistical model for these effects assumes that each interviewer $k$ biases the responses they collect slightly. The response to question $Y$ collected from respondent $i$ is a function of the true response and the interviewer level error term, $\delta_{k}$.
$Y^{measured}_{i,k}$ $=f(Y^{true}_{i},\delta_{k})$
$\delta_{k}$ $\sim N(0,\Delta)$
When a survey uses only a few interviewers, the individual effects can introduce bias into estimates. When there are many interviewers, the individual effects cancel out. For this reason, it is best practice in surveys to use many interviewers and make sure no interviewer dominates the data set (Kish, 1962; Groves, 1989).
Questions eliciting respondents’ opinions have larger interview effects than those about factual matters (that is, a larger $\Delta$). The explanation is the same as that cited above to explain why opinion questions are more susceptible to context effects: respondents often form opinions when asked for them, rather than drawing on pre-formed opinions. When they do so, they are impacted by the interviewer’s appearance, tone, demographics, personality, etc. (Schnell & Kreuter, 2005; West & Blom, 2016). However, interviewer effects also arise with factual and even demographic questions (Hansen et al., 1960).
4.2 Labeler Effects
In a similar way, each labeler $k$ may have a small effect on the labels they assign, $L^{collected}$, which we might call $\lambda_{k}$.
$L^{collected}_{i,k}$ $=f(L^{true}_{i},\lambda_{k})$
$\lambda_{k}$ $\sim N(0,\Lambda)$
Although the ML literature has not conceptualized labeler effects in this way, we find evidence in the literature that these $\lambda$ values are non-zero and correlate with labeler characteristics. For example, labeler age and education level influence whether they perceive comments on Wikipedia entries as attacks (Al Kuwatly et al., 2020). Conservative labelers are less likely to flag anti-Black language (Sap et al., 2022). [^3] To our knowledge, most evidence of labeler effects comes from the NLP literature. However, our experience with interviewer effects across all types of questions leads us to believe that labeler effects also occur in image labeling, human feedback for reinforcement learning, and all types of label collection. As with surveys, the labeler effects are likely smaller for some tasks than others. More objective tasks that most labeler agree on may have small effects. Tasks like hate speech annotation probably have larger $\lambda$ values. These labeler effects also impact models: training a model on subsets of labelers with shared characteristics leads to different predictions (Al Kuwatly et al., 2020; Perikleous et al., 2022).
Just as with interviewers, it is preferable to have many labelers each with a reasonable number of labels rather than a few labelers with many labels. The argument is the same: When many labelers contribute to a data set, their individual effects cancel out and do not result in systematic bias. As discussed above, studies may find it impractical to have many labelers.
Although many consider the $\lambda$s noise, as we do in surveys, others have argued that they are in fact information and can improve model performance. Embracing label variation may make models more trustworthy (Aroyo & Welty, 2015; Plank, 2022).
4.3 Who Works as Labelers
Training models on labels that reflect the views of the population will increase the inclusiveness and trustworthiness of models. However, our experience with survey data collection leads us to worry about how well labelers represent the population.
@@ 4.3 Who Works as Labelers
Large-scale annotation tasks, such as the reCAPTCHA tests, may collect labels from a representative slice of the population. However, the crowdworkers used in many label collection tasks are members of large panels such as Appen, Upwork, Scale AI, Prolific, or MTurk and do not reflect the US or world population. The workers who labeled data to fine-tune InstructGPT were members of Scale AI and Upwork; 22% were Filipino and 22% were Bangladeshi (Ouyang et al., 2022, Appendix B3). MTurk members are younger, lower income and less likely to live in the South than the US population (Berinsky et al., 2012).
@@
In surveys, selection bias [^4] In the survey literature, the preferred terms is nonresponse bias. We use the term selection bias here because it is more general and more suited to the labeling task. arises when the propensity to take part in a survey is correlated with the characteristics measured in a survey. Return to the example in Section 2, a survey to estimate the proportion of US adults without health insurance. If we distributed the survey invitations in doctors’ waiting rooms, the sample proportion would overestimate the population proportion. Those in waiting rooms are more likely to receive the survey invitation and are also more likely to have health insurance. The propensity to take part is correlated with what the survey measures, leading to selection bias in the estimate of health insurance coverage.
@@
Just as in surveys, selection bias arises in training data labels if the propensity to engage in the labeling task is correlated with the propensity to assign a given label. We expect that this correlation is non-zero for many tasks, for two reasons. The first reason we expect a correlation is that labeler characteristics likely influence both propensities, as shown in Figure 2. On the right side of the figure, labelers’ characteristics influence their decision to engage in labeling work, as discussed earlier in this subsection. On the left side of the figure, labelers’ characteristics influence the labels they assign, as discussed in the previous subsection on Labeler Effects (Section 4.2). If the characteristics that influence one side of the figure are the same as, or correlate with, the characteristics that influence on the other side of the figure, the two propensities at the bottom will be correlated, leading to selection bias.
The second reason we expect a correlation is that labeling tasks are largely opt-in and labelers can see the type of work they will be involved in before deciding to participate. For example, someone with strong views on bicycle safety may be more likely to see bike lane violations where others do not and, for that reason, be more likely to choose to label a data set of images of potential bike lane violations.
Figure 2: Labeler characteristics impact propensity to assign a given label and propensity to participate in labeling
@@ 4.4 Mitigation Measures and Future Research
To combat selection bias in labeling tasks, we need to break the correlation between the propensity to assign a given label and the propensity to participate in labeling. One way to do so is to remove the correlation between labeler characteristics and the propensity to participate in labeling by diversifying the labeler pool, collecting data from labelers with different motivations and characteristics. This approach is central to surveys: we solicit responses from a random sample of the population through appeals to public service, tokens of appreciation, and multiple reminders (see, for example, Groves & Couper, 1998). If we can collect responses from labelers who represent the population, selection bias is not a problem. Unfortunately, it is not clear that this approach will work with label collection: many people are not interested in labeling data for AI models. And some tasks require expert knowledge, such as reading radiographs or annotating low-resource languages; for these tasks, representative samples are likely not possible. (Surveys also find it increasingly difficult to recruit representative samples of respondents (de Leeuw E., 2018; Williams & Brick, 2018)).
@@
Another way to break the correlation in Figure 2 is through statistical adjustment. The survey literature contains many statistical methods to match the characteristics of the respondents to the population and thus reduce selection bias (Bethlehem et al., 2011, for example). For example, surveys in many countries collect more responses from women than men; we use weights to equalize the contribution of women’s and men’s responses on the final estimate. Future work could test the usefulness of these weighting approaches for improving machine learning models. [^5] The tendency for some crowdworkers to give false answers to demographic questions to protect their privacy, reported by Huang et al. (2023), will complicate any weighting approach. However, these statistical adjustments can work only if we know and measure the labeler characteristics that drive the relationships in Figure 2. Thus we need a better understanding of what motivates people to work as labelers and what types of tasks are vulnerable to labeler effects (the two paths in Figure 2), which points to another role for social science to play in improving label quality.
@@ 5 Transparency in Label Collection
The discussion above suggests that labels are sensitive to how studies design the labeling task and recruit, train, and compensate labelers, in ways not recognized in the AI/ML literature. For this reason, we call for more transparency and documentation in how labels are collected when new data sets or models are released.
The survey industry in the U.S. has embraced transparency in recent years. The American Association for Public Opinion Research launched the Transparency Initiative in 2014. Survey data collection firms that are accepted as members agree to disclose details about how survey data were collected when requested, such as sponsor, question wording, and weighting adjustments. [^6] See https://aapor.org/wp-content/uploads/2022/11/TI-Attachment-C.pdf for details. Polling companies that are members of the Transparency Initiative outperform those who are not, suggesting that willingness to disclose methods is a proxy for the quality of a firm’s data collection methods (Silver, 2023). The U.S. federal statistical agencies recently commissioned an expert report on how they can improve transparency as well (National Academies of Sciences, Engineering, and Medicine, 2022), another sign that transparency in data collection has positive externalities.
Several researchers have similarly called for transparency when releasing benchmark data sets for AI and ML models (Bender & Friedman, 2018; Hutchinson et al., 2020; Gebru et al., 2021; Chmielinski et al., 2022). However, none of these standards request the kind of data collection process information required by the Transparency Initiative. The absence of more detailed documentation of data collectinon methods may be behind the recent difficulties in replicating benchmark data sets (Recht et al., 2019). We commend Nie et al. (2020); Ouyang et al. (2022) as particularly good examples of transparency in label collection, with the latter including screenshots of the label collection instrument (Figure 1). Prabhakaran et al. (2021); Geiger et al. (2020); Baan et al. (2024) have also called for better documentation of the label collection process.
@@ 6 Outlook
Collecting data from respondents or labelers is more difficult than the AI/ML literature has recognized. Opinion questions are particularly difficult, because opinions can be shaped by wording, order, and other context effects and by the characteristics of the interviewer or the labeler. We believe that the type of human feedback needed to ensure that future models are trustworthy will likely resemble opinion collection in surveys, as in Figure 1. For this reason, greater cooperation and knowledge sharing between the AI/ ML and survey methods fields will be crucial to ensuring that the next generation of even more powerful models is human-centric, trustworthy, and fair.
@@
Although some hope to replace human labelers with models-as-labelers (Pangakis et al., 2023), the data collection challenges described in this paper will not go away. First, because models are trained on data collected from humans, they can display many of the response biases described in this paper (Dominguez-Olmedo et al., 2023; Tjuatja et al., 2023). Second, models trained on data labeled by models exhibit unusual behavior, called Model Autophagy Disorder (Alemohammad et al., 2023). Third, for the foreseeable future, we will still want humans to create the labels that teach models how to be fair and safe. Thus, the most difficult and most important labeling must still be done by humans.
As is often the case when working across disciplines, as this paper does, language and cultural differences get in the way of information sharing. The survey methodology literature tends to focus on uncovering bias and variance and attributing it to a cause (interviewers, question wording, survey mode). AI researchers, on the other hand, often understandably focus on solutions in the form of python packages or github repos. These cultural differences between the fields can leave researchers unable to present at each other’s conferences or even have productive conversations. Nevertheless, we hope that this position paper has made the case that theory-driven social science can help AI researchers develop insights and tools that improve the quality of their data and their models. And solution-oriented thinking could also make survey data better. Perhaps future joint workshops could help the two fields share insights.
@@
We applaud the growing interest in data-centric AI, which focuses on improving models by improving the data on which they rely rather than using the latest algorithm. We recommend that researchers developing AI models be as careful with how their data are collected as they are with their models. We also suggest reaching out to applied statistics departments and social science research groups which often focus on the science of data and may be willing to do the data work.
Acknowledgements

### v2

@@ 3.3 Mitigation Measures and Future Research
The survey methods literature suggests several approaches to minimize the effects discussed in Section 3.2.
Randomization of Observations
To address order effects, label collectors can randomize the order of observation shown to labelers. This approach does not eliminate order effect but it ensures that no one ordering impacts all annotators in the same way. Random ordering is incompatible with active learning techniques, however.
@@
Many surveys spend weeks or even months drafting, testing, and revising questions and response options to arrive at language that is understood similarly by most members of the population, a process called cognitive interviewing (Willis, 2004). They then launch the survey with a small group of respondents to assess response rates, don’t know rates, and response times. Such testing could improve the instructions and prompts given to labelers.
@@
Many surveys capture process data, called paradata, during the survey, such as the time spent on each screen, the device used, even mouse movements (Kreuter, 2013; Horwitz et al., 2017). Paradata can help identify satisficing respondents and low quality data (Kreuter, 2013) and may do the same in label collection. However, collecting such data may raise additional privacy and ethics concerns (see Couper & Singer, 2012; Kunz et al., 2020; Henninger et al., 2023).
@@
Label collection instruments could experiment with prompts to encourage labelers not to engage in satisficing. Respondents who pick many “don’t know” answers or repeatedly choose the same response option could receive reminders about the importance of the task. Those who click through screens quickly could receive prompts to slow down and read carefully. In surveys, feedback on speeding successfully slowed respondents and did not lead to early terminations of the survey (Conrad et al., 2017).
@@
In surveys, instructed response items, such as “Choose yellow below” can help identify respondents who speed or provide low quality responses (Gummer et al., 2018; Berinsky et al., 2024). We have not seen these questions used in labeling tasks. However, some tasks embed observations with known labels to try to catch annotators who do not understand the task. In NLP tasks, it is common to qualify only workers who pass an initial quiz or perform well on inserted test observations (Nangia et al., 2021). These test observations could also catch annotators who satisfice (see Nie et al., 2020, for an application).
@@
We recommend future research to more comprehensively test these and related approaches.
4 Need for Diversity
@@
Large-scale annotation tasks, such as the reCAPTCHA tests, may collect labels from a broad spectrum of the population. However, the crowdworkers used in many label collection tasks are members of large crowdworker panels such as Appen, Upwork, Scale AI, Prolific, or MTurk and do not reflect the U.S. or world population. Smart et al. (2024) note that labelers tend to be from the Global South, while the models they help train benefit the educated population in the Global North. The workers who labeled data to fine-tune InstructGPT were 22% Filipino and 22% Bangladeshi (Ouyang et al., 2022, Appendix B3). MTurk members are younger, lower income and less likely to live in the South than the U.S. population (Berinsky et al., 2012).
The unique characteristics of the labelers lead us to worry that the data they provide may not represent the views of the population that will use or be affected by the models. [^2] We acknowledge that this section glosses over what we mean by “population.” Is it the population that regularly uses the models? The population impacted by the models? We leave this important discussion to later work. Issues of representativeness are enormously important to surveys, which explicitly aim to make statements about the entire population. For this reason, the survey methods literature has much to contribute on this topic.
4.1 Selection Bias
@@
Selection bias occurs when those involved in providing data have different characteristics than the population. In surveys, selection bias [^3] In the survey literature, the preferred terms is nonresponse bias. We use the term selection bias here because it is more general and more suited to the labeling task. arises when the propensity to take part in a survey is correlated with the characteristics measured in a survey. Let us return to the example in Section 2, a survey to estimate the proportion of U.S. adults without health insurance. If we distributed the survey invitations in doctors’ waiting rooms, the sample proportion would overestimate the population proportion. Those in waiting rooms are more likely to receive the survey invitation and are also more likely to have health insurance. The propensity to take part is correlated with what the survey measures, leading to selection bias in the estimate of health insurance coverage.
@@
Just as in surveys, selection bias arises in training data labels if the propensity to engage in the labeling task is correlated with the propensity to assign a given label. We expect that this correlation is non-zero for many tasks, because labeler characteristics likely influence both propensities, as shown in Figure 3.
Figure 3: Labeler characteristics induce correlation between propensities (adapted from Groves, 2006)
On the left side of Figure 3, labelers’ characteristics influence their decision to engage in labeling work. As noted above, labelers tend to have different attributes than the general population (Gray & Suri, 2019; Smart et al., 2024). On the right side of Figure 3, labelers’ characteristics influence the labels they assign. The literature provides evidence for this association as well: labeler age and education level impact whether they perceive comments on Wikipedia entries as attacks (Al Kuwatly et al., 2020); conservative labelers are less likely to flag anti-Black language (Sap et al., 2022); labelers in the U.S. are more likely to see a bird in ambiguous images than those in India (Parrish et al., 2024). We hypothesize that such effects are more widespread than the literature suggests, because many studies do not collect annotator characteristics and thus cannot detect their impact on the labels (Kirk et al., 2023).
If the characteristics that influence one side of Figure 3 are the same as, or correlate with, the characteristics that influence the other side of the figure, the two propensities at the bottom will be correlated, leading to selection bias. Consider a labeler who is a frequent biker, annoyed by cars that park in the bike lanes in their city. They may be more likely to agree to label a data set of potential bike lane violations (left side) and also more likely to see violations where others do not (right side).
Although the goal of label data collection is not to make population estimates, selection bias is nevertheless a risk to training data and model development. In the early days of machine learning, when developers trained model to recognize written numbers or tell cats from dogs, perhaps labeler characteristics mattered less. The association on the right side of Figure 3 may be weak or absent with more objective tasks (though see Aroyo & Welty, 2015, for counterarguments).
However, the labeling tasks still performed by humans today often involve more difficult and more opinion-based work. People may legitimately disagree about whether a given statement is toxic or offensive, for example. Reinforcement learning with human feedback (RLHF) in particular may be more exposed to selection bias. The labels collected for reinforcement learning, like those in Figure 1, are inherently opinion-based. As discussed in Section 2, opinion questions in surveys are more susceptible to context effects than factual questions. We suspect that opinion labels are also more impacted by labeler characteristics than are more objective labels. RLHF aligns models to the judgments of labelers. If selection bias is present in the data, those judgments do not match the interests of the public. For example, if those who participate in labeling are also less likely to judge text as toxic, then the model trained on their data will also see less toxicity.
We see evidence of the impact of selection bias on models in two studies that trained models on different sets of labelers. Both showed that models make different predictions when trained on labels from, for example, female versus male labelers, or Asian versus white labelers (Al Kuwatly et al., 2020; Perikleous et al., 2022).
@@ 4.2 Mitigation Measures and Future Research
To combat selection bias in labeling tasks, we need to break the correlation between the propensity to assign a given label and the propensity to participate in labeling. We consider three methods.
Left Side
@@
We could try to remove the correlation between labeler characteristics and the propensity to participate in labeling (the left side of Figure 3) by diversifying the labeler pool, collecting data from labelers with different motivations and characteristics. This approach is central to surveys: we solicit responses from a random sample of the population through appeals to public service, tokens of appreciation, and multiple reminders (see, for example, Groves & Couper, 1998). If we can collect responses from a random sample of population members, selection bias is not a problem. Unfortunately, it is not clear that this approach will work with label collection: many people are not interested in labeling data for AI models. Surveys also find it increasingly difficult to recruit representative samples of respondents (de Leeuw et al., 2018; Williams & Brick, 2018).
Right Side
Another way to reduce selection bias is to break the correlation on the right side of Figure 3, which would mean removing (or, more reasonably, reducing) the influence of labeler characteristics on the labels that they assign. More diverse examples in the instructions, use of test observations, feedback to labelers, and training in implicit bias might help labelers label more uniformly. (The literature on coding in qualitative studies takes a similar approach (see, for example, Hak & Bernts, 1996).) Even if these interventions do work to reduce the impact of labelers’ characteristics on the labels they provide, however, they are expensive and do not scale well.
Interestingly, reducing selection bias by removing the correlation on the right side of the figure is not of interest in surveys, which aim to capture the diversity of respondents’ behavior, opinions, and judgments. Recent research in NLP has similarly found that capturing the diversity of labels across labelers can improve models (Basile et al., 2021; Sap et al., 2022). Aroyo & Welty (2015) and Plank (2022) have argued that such human label variation is in fact information and can improve model performance and trustworthiness.
Weighting
A third method to address selection bias is statistical adjustment. If we condition on the labeler characteristics in Figure 3, we can remove (or reduce) the induced correlation between the two propensities. Like the right side approach, this method involves embracing labeler subjectivity and uses weights to get the balance of those characteristics right.
@@
The survey literature contains many statistical methods to match the characteristics of the respondents to the population and thus reduce selection bias (Bethlehem et al., 2011, for example). For example, surveys in many countries collect more responses from women than men; we use weights to ensure that the contribution of women’s and men’s responses on the final estimate matches their shares in the population. Future work could test the usefulness of these weighting approaches for improving machine learning models. [^4] The tendency for some crowdworkers to give false answers to demographic questions to protect their privacy, reported by Huang et al. (2023), will complicate any weighting approach. Misreporting of demographics also causes problems in surveys (Pew Research Center, 2024).
However, these statistical adjustments can work only if we capture the labeler characteristics that drive the relationships in Figure 3. Thus we need a better understanding of what motivates people to work as labelers and what types of tasks are vulnerable to selection bias, which points to another role for social science to play in improving label quality.
@@ 5 Transparency in Label Collection
The discussion above suggests that labels are sensitive to how studies design the labeling task and recruit labelers, in ways often not recognized in the AI/ML literature. For this reason, we call for more transparency and documentation in how labels are collected when new data sets or models are released.
The survey industry in the U.S. has embraced transparency in recent years. The American Association for Public Opinion Research launched the Transparency Initiative in 2014. Member firms agree to disclose details about how survey data were collected, such as question and response option wording and order, respondent recruitment protocols, and weighting adjustments. [^5] See https://aapor.org/wp-content/uploads/2022/11/TI-Attachment-C.pdf for details. Polling companies that are members of the Transparency Initiative outperform those who are not (Silver, 2023), suggesting that a firm’s willingness to disclose its data collection methods is a proxy for the quality of its estimates. The U.S. federal statistical agencies recently commissioned an expert report on transparency, in an effort to increase trust in federal data (National Academies of Sciences, Engineering, and Medicine, 2022).
Several researchers have similarly called for transparency when releasing benchmark data sets or models (Bender & Friedman, 2018; Mitchell et al., 2019; Hutchinson et al., 2020; Gebru et al., 2021; Chmielinski et al., 2022). We join these calls and recommend releasing the labeling instructions or guidelines including examples and test questions, the wording of the prompts, information about the labelers, and whether social scientists or domain experts were involved in labeling or consulted on the labeling process. Prabhakaran et al. (2021); Geiger et al. (2020); Ulmer et al. (2022); Baan et al. (2024) have also called for better documentation of the label collection process. We commend Nie et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Bai et al. (2022); Stiennon et al. (2022) as particularly good examples of transparency in label collection, with some including screenshots of the label collection instrument.
Without detailed documentation of data collection methods, researchers will not be able to test many of the hypotheses given above, such as those about wording, task order, “don’t know” options, and the impact of labeler characteristics. We also suspect that lack of documentation explains difficulties replicating benchmark data sets (Recht et al., 2019).
@@ 6 Outlook
Collecting data from labelers is more difficult than the AI/ML literature has recognized. Ambiguous and opinion tasks are particularly challenging, because labelers’ responses can be shaped by wording, order, and other context effects as well as by the characteristics of the labeler. We believe that the type of human feedback needed to align future models will resemble opinion collection in surveys. For this reason, greater cooperation and knowledge sharing between the AI/ML and survey methods fields will be crucial to ensuring that the next generation of even more powerful models is human-centric, trustworthy, and fair.
@@
Although some hope to replace human labelers with models-as-labelers (see, for example, Pangakis et al., 2023; Gilardi et al., 2023; Törnberg, 2023), the data collection challenges described in this paper will remain. Because models are trained on data collected from humans, as labelers, they can display many of the response biases described in this paper (Dominguez-Olmedo et al., 2023; Tjuatja et al., 2023). In addition, models trained on data labeled by models exhibit unusual behavior, called model collapse or model autophagy disorder (Alemohammad et al., 2023; Gerstgrasser et al., 2024; Peterson, 2024). Models trained on data labeled by models may also exhibit lower performance than those trained on data labeled by humans (see, for example, Plaza-del Arco et al., 2024). For the foreseeable future, we will still want humans to create the labels that teach models how to be accurate, fair and safe. Thus, the most important labeling must still be done by humans and is exactly the type of data that is most challenging to collect.
@@
We applaud the growing interest in data-centric AI, which focuses on improving models by improving the data they rely on. We recommend that researchers developing AI models be as careful with how their data are collected as they are with their models. We also suggest reaching out to applied statistics departments and social science research groups which often focus on the science of data and may be willing to do the data work.
As is often the case when working across disciplines, as this paper does, language and cultural differences complicate information sharing. The survey methodology literature tends to focus on detecting and measuring bias and variance and attributing it to a cause (interviewers, question wording, survey mode). AI researchers, on the other hand, often understandably focus on solutions, in the form of algorithms distributed through python packages or github repos. These cultural differences between the fields can leave researchers unable to present at each other’s conferences or even have productive conversations. Nevertheless, we hope that this position paper has made the case that theory-driven social science can help AI researchers develop insights and tools that improve the quality of their data and their models. [^6] We acknowledge that practitioners in AI and ML may be more aware of findings from survey research than the literature suggests. (And solution-oriented thinking could also make survey data better.) Perhaps future joint workshops could help the two fields share insights.
Impact Statement
This position paper uses results from the field of survey methodology to derive hypotheses about the drivers of data quality in training data collection. We propose several areas for future research and provide concrete ideas to improve the quality of labels collected to train, fine-tune, reinforce, and evaluate AI and ML models. By improving the labels, these ideas will have positive effects on model performance and alignment.
Acknowledgments
In writing this paper, we benefited from discussions with Jacob Beck and Bolei Ma of LMU and Rob Chew of RTI International. Fiona Draxler provided valuable comments on an early draft. We would also like to thank the anonymous reviewers for their feedback. BP is supported by the European Research Council (ERC) grant agreements No. 101043235. This research is funded in part by the Bavarian Research Institute for Digital Transformation (bidt) project KLIMA-MEMES (BP) and GREEN DIA (FK), and the German Federal Ministry of Education and Research (BMBF) project KODAQS (16DKZ2019C). The authors are responsible for the content of this publication.